OpenStudy (anonymous):

Find the value of $5000 compounded quarterly at 8% for 10 years

OpenStudy (anonymous):

A=p( 1+r)^t so plug in numbers, A= 5000( 1 + .08)^10 and solve

OpenStudy (anonymous):

i'm sorry correction

OpenStudy (anonymous):

The formula would be A= p (1+r/n)^nt so.... A= 5000( 1 + .08/4)^4+10

OpenStudy (anonymous):

No the answer would be 11040.20. i'll do the work to show you

OpenStudy (anonymous):

A=5000(1+.08/4)^4*10= 5000(1.02)^40=5000 (2.208039664)= 11040.198

OpenStudy (anonymous):

the interest is compounded quarterly at 8% which is .08 and it's divided by 4 because it's quarterly. then the quarters are multiplied times the year which stands for nt and that gives you the 40.

OpenStudy (anonymous):

I just rounded on the answers. but hope that helps.